OpenStudy (anonymous):

Please help, need it so much! [9.06] Jacob kicks a soccer ball off the ground and in the air with an initial velocity of 33 feet per second. Using the formula H(t) = −16t2 + vt + s, what is the maximum height the soccer ball reaches? 15.1 feet 16.5 feet 17.0 feet 18.2 feet

OpenStudy (anonymous):

I don't kno if I am rite, but is it the first one?

OpenStudy (anonymous):

well, when you want the maximum or minimum of any function you differentiate and put it equal to zero and find out what that variable was, but i think its a different matter here, give me a sec.

OpenStudy (anonymous):

Ok, I will, thx. :)

OpenStudy (anonymous):

what is the formula?

OpenStudy (anonymous):

H(t) = −16t2 + vt + s

OpenStudy (anonymous):

as in what is H(t)? height at given time?

OpenStudy (anonymous):

Yes, I think so. :)

OpenStudy (anonymous):

It is a height/ time formula. :)

OpenStudy (anonymous):

what is the variable s? is this the full question you dont mind me asking,

OpenStudy (anonymous):

Okay, were you given s? or is s=0?

OpenStudy (anonymous):

I think it is seconds. :) Yes, it is the full ?, why?

OpenStudy (anonymous):

Wait, when you answer, I will come back later and solve it. :)

OpenStudy (anonymous):

s is distance, no? ok, i will try to answer your question but it will require your judgement, if i did it wrong or not.. H(t)=-16t^2+vt+s okay, if we differentiate (wrt t) and set to 0 then we can find the maximum, right? so... d H(t) = dt -32t+v+(d/dt)[s] ---> differentiate s with respect to t we know that if we differentiate s wrt t we get velocity, right? sooooo.. the equation becomes -32t+v+v which is -32t+2v now we set it to zero and we know v, which is 33 feet per second, so, -32t+66=0 solve for t which is t=66/32 now this is the time which the ball is at its maximum, subbing this in the function H(t)...

OpenStudy (anonymous):

this is how far i got, i hope if someone else answers this question maybe they will carry on from there How did you get your answer to be the first one?

OpenStudy (anonymous):

Hmm...I don't know actually, I basically just guessed. :)

OpenStudy (anonymous):

I got it, s is distance, no? s=0 as its being hit from the ground! H(t)=-16t^2+vt+s okay, if we differentiate (wrt t) and set to 0 then we can find the maximum, right? so... d H(t) = dt -32t+v sooooo.. the equation becomes -32t+v now we set it to zero and we know v, which is 33 feet per second, so, -32t+33=0 solve for t which is t=33/32 now this is the time which the ball is at its maximum, subbing this in the function H(t)... we get H[33/32]= -16*(33/32)^2 +(33/32)*33 + 0 which is 17.0156, im still not too sure as it seems sooooooo trivial!!!

OpenStudy (anonymous):

Ok, thx so much! This is really hard stuff, lol!